Determine how many solutions exist.
Use either elimination or substitution to find the solutions (if any).
Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection.
y = 2x + 3 and y = -x - 4
I have found that there is one solution. Here's what I have so far:
First I subtracted the first equation from the second equation.
y = -x - 4
-y =-2x - 3 --> 0 =-3x - 7 I add 7 to both sides to get: 7 = -3x
Then I divide both sides by -3 to get x = -7/3 or x = -2.333
Then I substituted -2.333 in for x in the first equation to find y which came out to -1.666 (when x is -2.333). I came up with intercept points reduced to the hundreths as (-2.33,-1.67) to get the intercept point. Thanks.

Answer by mananth(16949)

You are almost there.
-2 x + 1 y = 3 .............1
1 x + 1 y = -4 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
2 x -1 y = -3
1 x 1 y = -4
Add the two equations
3 x = -7
/ 3
x = -2.33 OR -2 1/ 3
plug value of x in (1)
-2 x + 1 y = 3
4.67 + 1 y = 3
1 y = 3 -4.67
1 y = -1.67
y = -1.67
equation1
when x= 0 y= 3 ( 0 , 3 ) 3
when y = 0 x= -1.5 ( -1.5 , 0 )
Equation 2
when x= 0 y= -4 ( 0 , -4 ) -4
when y = 0 x= -4 ( -4 , 0 )